# Ideal Triangulations and Once-Punctured Surface Bundles

### v1

@@ 3.3. Short Inflations
(between band tetrahedra) Let $b_{i},b_{j}$ be two band tetrahedra identified at face $\sigma$ containing a trapezoidal face of $\mathcal{C}_{X}$. Each vertex of the associated bi-pyramid, $P$, gets identified to the sole vertex of $\mathcal{T}$ and a single edge in $P$ is the lift of some edge, $e$ in the material boundary of $\mathcal{T}$. Consider the lifts of the intersection of the boundary-linking surface $B$ with $P$. After the 2-3 move at $\sigma$, suppose $\widetilde{\Delta}_{\tilde{e}}$ is the unique tetrahedra in $P$ containing an edge, $\tilde{e}$, in the lift of $e$. The associated normal discs of $B^{\prime}$ in $\widetilde{\Delta}_{\tilde{e}}$ are then a normal quadrilateral separating $\tilde{e}$ and two normal triangles separating the two vertices not in $\tilde{e}_{k}$. The associated normal discs in the remaining two tetrahedra $\widetilde{\Delta}_{2},\widetilde{\Delta}_{3}$ of $P$ are one of each of the four normal classes of normal triangle discs in each respective tetrahedra. As $\mathcal{C}_{X}$ is simple, by Remarks 3.3 and 3.2 the cell-decomposition $\mathcal{C}^{\prime}_{X}$ induced on $X$ by $\mathcal{T}^{\prime}$ is simple. Further the preimage of a component $\mathbb{P}^{\prime}_{i}$ of $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ is a component $\mathbb{P}_{j}$ of $\mathbb{P}(\mathcal{C}_{X})$. Thus $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ is a trivial combinatorial product. By Theorem 3.1, $\mathcal{T}^{\prime}$ is an inflation triangulation of $M$. The number of band tetrahedra $\beta^{\prime}=\beta-1$, the number of crossing tetrahedra $\gamma^{\prime}=\gamma$; hence $len(\mathcal{T}^{\prime})=len(\mathcal{T})-1$.
@@
There exists an algorithm [5, 21, 34] to construct an ideal triangulation $\mathcal{T}^{*}$ of $M$. Let $\mathcal{T}_{\xi}$ be an inflation of $\mathcal{T}^{*}$ constructed as in [23] by way of the frame $\xi$. By Case 1 of Proposition 3.4, we may form an inflation triangulation $\mathcal{T}^{\prime}$ having no crossing tetrahedra. As $\xi$ is a frame and $\partial M$ is a torus, $\xi$ must consist of exactly two or three branches.
@@
We will now establish a relationship between normal surfaces with boundary in a short inflation, $\mathcal{T}$ and spun-normal surfaces in the ideal triangulation obtained by crushing $\mathcal{T}$ along $\partial M$. First, we note as $\mathcal{T}$ is a minimal-vertex triangulation of $M$, $\partial M$ has an induced triangulation given by the two unglued faces of $\mathcal{T}$. For a normal surface $S$ with nonempty boundary, $\partial S$ has an induced cell-decomposition consisting of normal arcs in the material boundary of $\mathcal{T}$. If $S$ is incompressible and $\partial$-incompressible, each component of $\partial S$ is non-trivial in $\pi_{1}(\partial M)$. The topology of $\partial M$ then imposes a restriction on the normal arcs of $\partial S$.
@@
For a short inflation, the normal discs of a surface $S$ in the band tetrahedra together with the normal arcs of $\partial S$ full determine the normal discs of $S$ in the branch-point tetrahedra. The result then follows immediately from the restriction that no component of $\partial S$ bounds a disc in the material boundary of $\mathcal{T}$ together with Lemma 3.8.
We now come to the main tool of this work: the construction of a short inflation in which a surface is compatible. It is compatibility that will ensure the existence of a spun-normal surface upon crushing the boundary of $\mathcal{T}$.
@@
The material boundary of $\mathcal{T}$ is given by the two unglued faces of the branch-point tetrahedra $p_{1}$ and $p_{2}$. There are three edges in the material boundary of $\mathcal{T}$, one with order 2 and the others having order 3. Orient and label the two order 3 edges $A,B$ as shown in Figure 5. The fundamental group $\pi_{1}(\partial M)$ is generated by the closed curves $A$ and $B$. The boundary curve $\partial S$ can be given an orientation so that $\partial S=nA+mB$, with $\gcd(m,n)=1$. As in Lemma 3.8 and Figure 4, we refer to the six normal arc types in the material boundary of $\mathcal{T}$ as $\alpha,\alpha^{\prime},\beta,\beta^{\prime},\gamma,\gamma^{\prime}$. Let $x_{\alpha},y_{\alpha}$ count the number of normal triangles and normal quadrilaterals respectively in the cell-decomposition of $S$ with normal arc type $\alpha$; similarly, for the other normal arc types define such $x_{\alpha^{\prime}},y_{\alpha^{\prime}}\dots$ Note that as $S$ is $\partial$-irreducible, it cannot be the case that both $x_{\gamma}$ and $x_{\gamma^{\prime}}$ are positive. The branch-point tetrahedra have a single face, $\sigma$, in common with the lifts $\tilde{p}_{1},\tilde{p}_{2}$ meeting at $\sigma$ forming a square pyramid $P$ with base the unglued faces of $\tilde{p}_{1},\tilde{p}_{2}$. Up to reflective symmetry of the material boundary of $\mathcal{T}$ and relabeling the edges and possibly performing a 2-2 move on $P$, it can be assumed $0\leq m\leq n$. Thus we need only to prove the result when $\partial S$ has normal arc types of $\alpha,\alpha^{\prime},\gamma,\gamma^{\prime}$. That is, $x_{\beta},x_{\beta^{\prime}},y_{\beta},y_{\beta^{\prime}}=0$.
@@
Therefore, if $S$ a connected, properly embedded, incompressible, $\partial$-incompressible surface in $M$ with $\partial S$ connected, the short inflation $\mathcal{T}$ can thus be altered through successive site-swapping moves and shortening moves into a short inflation in which $S$ is isotopic to a compatible normal surface.
@@
Let $M$ be an orientable, compact, irreducible, $\partial$-irreducible 3-manifold with non-empty boundary homeomorphic to a torus. Further, let $\mathcal{T}$ be a short inflation with $S$ a compatible normal surface. The ideal triangulation $\mathcal{T}^{*}$ obtained by crushing $\mathcal{T}$ along $\partial M$ has a spun-normal surface $S^{*}$ with normal quadrilaterals completely determined by the normal quadrilaterals of $S$ in $\mathcal{T}$, the lifts of which (relative to the crushing map) are contained in $\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$. Further $S^{*}$ represents the interior of a properly embedded surface in $\mathring{M}$ isotopic to $\mathring{S}$.
@@
Let the 3-manifold $M$, short inflation $\mathcal{T}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$, and compatible normal surface $S$ be given as above. Let $\mathcal{T}^{*}$ be the ideal triangulation obtained by crushing $\mathcal{T}$ along $\partial M$. Let $t=|\widetilde{\mathbf{\Delta}}|$. Recall that we may write $\widetilde{\mathbf{\Delta}}=\widetilde{\mathbf{\Delta}}^{*}\cup\{\widetilde{b}_{1},\widetilde{b}_{2}\}\cup\{\widetilde{p}_{1},\widetilde{p}_{2}\}$ where $\widetilde{b}_{i}$ are band tetrahedra and $\widetilde{p}_{i}$ are branch-point tetrahedra in the short inflation. Further, we have $\mathcal{T}^{*}=(\widetilde{\mathbf{\Delta}}^{*},\mathbf{\Phi}^{*})$ described by a one-to-one correspondence between the ideal tetrahedra of $\mathcal{T}^{*}$ and those in $\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$ and $\mathbf{\Phi}^{*}$ induced by the crushing map. Through this one-to-one correspondence we associate the normal quadrilateral type $d^{Q}_{i}$ in $\widetilde{\Delta}_{j}\in\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$ to the normal quadrilateral type $d^{Q*}_{i}$ in the ideal tetrahedron $\widetilde{\Delta}^{*}_{j}\subset\widetilde{\mathbf{\Delta}}^{*}$ of $\mathcal{T}^{*}$. From the vector $S_{Q}=(q_{1},q_{2},\dots,q_{3t})\in\mathbb{R}^{3t}_{\geq 0}$ representing $S$ in $\mathcal{T}$ with $d^{Q}_{i}$ representing a normal quadrilateral type in some $\widetilde{b}_{1},\widetilde{b}_{2},\widetilde{p}_{1},\widetilde{p}_{2}$ for $i>3t-12$. We can then construct a vector $v^{*}_{Q}=(q^{*}_{1},q^{*}_{2},\dots,q^{*}_{3t-12})\in\mathbb{R}^{3t-12}_{\geq 0}$ with $q_{i}^{*}=q_{i}$ for $121\leq i\leq 3t-$ 12. We wish to show that $v^{*}_{Q}$ represents an admissible spun-normal surface in $\mathcal{T}^{*}$.
@@
If for each tetrahedron $\Delta$ in $\mathcal{T}$ with at least one edge identified to some $e_{k}$ lifts to a tetrahedron $\widetilde{\Delta}\in\widetilde{\mathbf{\Delta}}^{*}$, then $\{e_{k}\}$ is in fact a singleton and the sense for every normal quadrilateral in band and branch-point tetrahedra is zero.
Three edges of a branch-point tetrahedron $\widetilde{p}_{i}$ are lifts of interior edges of $\mathcal{T}$; if one of these three edges is in $\{e_{k}\}$, then all three are included say $e_{1},e_{2},e_{3}$. For any normal quadrilateral type in $\widetilde{p}_{i}$ we have the sum of the senses given by $\sum_{k=1}^{3}\epsilon_{k}=0$.
In a band tetrahedron $b_{i}$ there is exactly one edge, say $\langle 01\rangle$ in the material boundary of $\mathcal{T}$. Let $d^{Q}_{1}$ represent the normal quadrilateral type that separates the edges $\langle 01\rangle$ and $\langle 23\rangle$; let $d^{Q}_{2},d^{Q}_{3}$ be the remaining two types. As $S$ is compatible, the number of normal quadrilaterals in $S$ of types $d^{Q}_{2},d^{Q}_{3}$ are $q_{2}=q_{3}=0$. The sign of a normal quadrilateral of type $d^{Q}_{1}$ with respect to the edge $\langle 23\rangle$ is $\delta_{1}=0$. If one of the (unoriented) edges $\langle 02\rangle,\langle 03\rangle,\langle 12\rangle,\langle 13\rangle$ of $\widetilde{b}_{i}$ are in $\{e_{k}\}$, then all for must be in $\{e_{k}\}$. Denote these edges by $e_{1},e_{2},e_{3},e_{4}$ respectively (they must all be distinct by the inflation construction). Then for each type $d^{Q}_{i}\in\{d^{Q}_{1},d^{Q}_{2},d^{Q}_{3}\}$ we have $\sum_{k}\epsilon_{k,i}q_{i}=0$. See Figure 7 for an illustration of the signs that contribute to these sense calculations.
@@
Let $B$ be the boundary linking surface in $\mathcal{T}$. For a normal quadrilateral of $B$ in a band tetrahedron $\widetilde{b}_{i}$ there are two opposite edges, $\varepsilon_{1},\varepsilon_{2}$ crush to an edge in the frame with which $\mathcal{T}$ was inflated (see Figure 2). Through normal isotopies $B$ and $S$ can be positioned such that the intersection $\Gamma$ of $S$ with $B$ has no arcs in the normal quadrilaterals of $B$ that are isotopic to $\varepsilon_{1},\varepsilon_{2}$. Crushing those cells of $\mathcal{C}_{X}$, the cell-decomposition of the complement of $B$ not containing the vertex of $\mathcal{T}$, which are not truncated-tetrahedra yields a cell-decomposition $\bar{\mathcal{T}}$ of $M$ consisting of truncated tetrahedra. By the same arguments of Lemma 3.4 in [6], crushing these cells maps normal discs of $S$ to a discs in $\bar{\mathcal{T}}$. As $S$ is compatible, the above Q-matching equations show that these discs glue together to form a surface $\bar{S}$ that is isotopic to $S$ in $M$. Specifically, crushing induces a cell-like map between $\bar{S}$ and $S\cap\mathcal{C}_{X}$. By [3, 39] $\bar{S}$ and $S\cap\mathcal{C}_{X}$ are homeomorphic. By the preceding arguments, $\bar{S}$ is a core of a spun-normal surface, the normal quadrilaterals of which determined by those normal quadrilaterals of $S$ in $\widetilde{\mathbf{\Delta}}^{*}$. Hence $\bar{S}$ is a core of $S^{*}$ completing the proof.
While in general for a band tetrahedron $\widetilde{b}_{i}$ the edge $\langle 23\rangle$ may be identified to the same edge class in $\mathcal{T}$ as one of (unoriented) $\langle 02\rangle,\langle 03\rangle,\langle 12\rangle,\langle 13\rangle$, the ideal triangulations used in Lemma 3.6 via way of [21] can have no faces which are 3-folds or dunce hats [25], these edge classes in this work must be distinct.
As incompatible normal quadrilaterals intersect the material boundary of $\mathcal{T}$, any closed normal surface must be compatible. This gives an alternate proof to a special case of Theorem 3.5 of [6].

### v2

@@ 3.3. Short Inflations
(between band tetrahedra) Let $b_{i},b_{j}$ be two band tetrahedra identified at face $\sigma$ containing a trapezoidal face of $\mathcal{C}_{X}$. Each vertex of the associated bi-pyramid $P$ gets identified to the sole vertex of $\mathcal{T}$ and a single edge in $P$ is the lift of some edge, $e$ in the material boundary of $\mathcal{T}$. Consider the lifts of the intersection of the boundary-linking surface $B$ with $P$. After the 2-3 move at $\sigma$, suppose $\widetilde{\Delta}_{\tilde{e}}$ is the unique tetrahedra in $P$ containing an edge, $\tilde{e}$, in the lift of $e$. The associated normal discs of $B^{\prime}$ in $\widetilde{\Delta}_{\tilde{e}}$ are then a normal quadrilateral separating $\tilde{e}$ and two normal triangles separating the two vertices not in $\tilde{e}_{k}$. The associated normal discs in the remaining two tetrahedra $\widetilde{\Delta}_{2},\widetilde{\Delta}_{3}$ of $P$ are one of each of the four normal classes of normal triangle discs in each respective tetrahedra. As $\mathcal{C}_{X}$ is simple, by Remarks 3.3 and 3.2 the cell-decomposition $\mathcal{C}^{\prime}_{X}$ induced on $X$ by $\mathcal{T}^{\prime}$ is simple. Further the preimage of a component $\mathbb{P}^{\prime}_{i}$ of $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ is a component $\mathbb{P}_{j}$ of $\mathbb{P}(\mathcal{C}_{X})$. Thus $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ is a trivial combinatorial product. By Theorem 3.1, $\mathcal{T}^{\prime}$ is an inflation triangulation of $M$. The number of band tetrahedra $\beta^{\prime}=\beta-1$, the number of crossing tetrahedra $\gamma^{\prime}=\gamma$; hence $len(\mathcal{T}^{\prime})=len(\mathcal{T})-1$.
@@
Let such a 3-manifold $M$ be given, and let $\mathcal{T}^{*}$ be an ideal triangulation of $M$. By [5, 21, 34] there exists an algorithm to construct an ideal triangulation of $M$. Let $\mathcal{T}_{\xi}$ be an inflation of $\mathcal{T}^{*}$ constructed as in [23] by way of the frame $\xi$. By Case 1 of Proposition 3.4, we may form an inflation triangulation $\mathcal{T}^{\prime}$ having no crossing tetrahedra. As $\xi$ is a frame and $\partial M$ is a torus, $\xi$ must consist of exactly two or three branches.
@@
We will now establish a relationship between normal surfaces with boundary in a short inflation $\mathcal{T}$ and spun-normal surfaces in the ideal triangulation obtained by crushing $\mathcal{T}$ along $\partial M$. First, we note as $\mathcal{T}$ is a minimal-vertex triangulation of $M$, $\partial M$ has an induced triangulation given by the two unglued faces of $\mathcal{T}$. For a normal surface $S$ with nonempty boundary, $\partial S$ has an induced cell-decomposition consisting of normal arcs in the material boundary of $\mathcal{T}$. If $S$ is incompressible and $\partial$-incompressible, then each component of $\partial S$ is non-trivial in $\pi_{1}(\partial M)$. The topology of $\partial M$ then imposes a restriction on the normal arcs of $\partial S$.
@@
For a short inflation, the normal discs of a surface $S$ in the band tetrahedra together with the normal arcs of $\partial S$ fully determine the normal discs of $S$ in the branch-point tetrahedra. The result then follows immediately from the restriction that no component of $\partial S$ bounds a disc in the material boundary of $\mathcal{T}$ together with Lemma 3.8.
We now come to the main technical results of this work: the construction of a short inflation in which a surface is compatible. It is compatibility that will ensure the existence of a spun-normal surface upon crushing the boundary of $\mathcal{T}$.
@@
The material boundary of $\mathcal{T}$ is given by the two unglued faces of the branch-point tetrahedra $p_{1}$ and $p_{2}$. There are three edges in the material boundary of $\mathcal{T}$, one with order 2 and the others having order 3. Orient and label the two order 3 edges $A,B$ as shown in Figure 6. The fundamental group $\pi_{1}(\partial M)$ is generated by the closed curves $A$ and $B$. The boundary curve $\partial S$ can be given an orientation so that $\partial S=nA+mB$, with $\gcd(m,n)=1$. As in Lemma 3.8 and Figure 5, we refer to the six normal arc types in the material boundary of $\mathcal{T}$ as $\alpha,\alpha^{\prime},\beta,\beta^{\prime},\gamma,\gamma^{\prime}$. Let $x_{\alpha},y_{\alpha}$ count the number of normal triangles and normal quadrilaterals respectively in the cell-decomposition of $S$ with normal arc type $\alpha$; similarly, for the other normal arc types define such $x_{\alpha^{\prime}},y_{\alpha^{\prime}}\dots$ Note that as $S$ is $\partial$-irreducible, it cannot be the case that both $x_{\gamma}$ and $x_{\gamma^{\prime}}$ are positive. The branch-point tetrahedra have a single face $\sigma$ in common with the lifts $\tilde{p}_{1},\tilde{p}_{2}$ meeting at $\sigma$ forming a square pyramid $P$ with base the unglued faces of $\tilde{p}_{1},\tilde{p}_{2}$. Up to reflective symmetry of the material boundary of $\mathcal{T}$ and relabeling the edges and possibly performing a 2-2 move on $P$, it can be assumed $0\leq m\leq n$. Thus we need only to prove the result when $\partial S$ has normal arc types of $\alpha,\alpha^{\prime},\gamma,\gamma^{\prime}$. That is, $x_{\beta},x_{\beta^{\prime}},y_{\beta},y_{\beta^{\prime}}=0$.
@@
Therefore, if $S$ a connected, properly embedded, incompressible, $\partial$-incompressible surface in $M$ with $\partial S$ connected, then the short inflation $\mathcal{T}$ can be altered through successive site-swapping moves and shortening moves into a short inflation in which $S$ is isotopic to a compatible normal surface.
@@
Let $M$ be an orientable, compact, irreducible, $\partial$-irreducible 3-manifold with non-empty connected boundary homeomorphic to a torus. Let $\mathcal{T}$ be a short inflation with $S$ a compatible normal surface. The ideal triangulation $\mathcal{T}^{*}$ obtained by crushing $\mathcal{T}$ along $\partial M$ has a spun-normal surface $S^{*}$ with normal quadrilaterals completely determined by the normal quadrilaterals of $S$ in $\mathcal{T}$, the lifts of which (relative to the crushing map) are contained in $\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$. Further, $S^{*}$ represents the interior of a properly embedded surface in $\mathring{M}$ isotopic to $\mathring{S}$.
@@
Let the 3-manifold $M$, short inflation $\mathcal{T}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$, and compatible normal surface $S$ be given as above. Let $\mathcal{T}^{*}$ be the ideal triangulation obtained by crushing $\mathcal{T}$ along $\partial M$. Let $t=|\widetilde{\mathbf{\Delta}}|$. Recall that we may write $\widetilde{\mathbf{\Delta}}=\widetilde{\mathbf{\Delta}}^{*}\sqcup\{\widetilde{b}_{1},\widetilde{b}_{2}\}\sqcup\{\widetilde{p}_{1},\widetilde{p}_{2}\}$ where $\widetilde{b}_{i}$ are band tetrahedra and $\widetilde{p}_{i}$ are branch-point tetrahedra in the short inflation. Further, we have $\mathcal{T}^{*}=(\widetilde{\mathbf{\Delta}}^{*},\mathbf{\Phi}^{*})$ described by a one-to-one correspondence between the ideal tetrahedra of $\mathcal{T}^{*}$ and those in $\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$ and $\mathbf{\Phi}^{*}$ induced by the crushing map. Through this one-to-one correspondence we associate the normal quadrilateral type $d^{Q}_{i}$ in $\widetilde{\Delta}_{j}\in\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$ to the normal quadrilateral type $d^{Q*}_{i}$ in the ideal tetrahedron $\widetilde{\Delta}^{*}_{j}\subset\widetilde{\mathbf{\Delta}}^{*}$ of $\mathcal{T}^{*}$. From the vector $S_{Q}=(q_{1},q_{2},\dots,q_{3t})\in\mathbb{R}^{3t}_{\geq 0}$ representing $S$ in $\mathcal{T}$ with $d^{Q}_{i}$ representing a normal quadrilateral type in some $\widetilde{b}_{1},\widetilde{b}_{2},\widetilde{p}_{1},\widetilde{p}_{2}$ for $i>3t-12$. We can then construct a vector $v^{*}_{Q}=(q^{*}_{1},q^{*}_{2},\dots,q^{*}_{3t-12})\in\mathbb{R}^{3t-12}_{\geq 0}$ with $q_{i}^{*}=q_{i}$ for $121\leq i\leq 3t-$ 12. We wish to show that $v^{*}_{Q}$ represents an admissible spun-normal surface in $\mathcal{T}^{*}$.
@@
If each tetrahedron $\Delta$ in $\mathcal{T}$ with at least one edge identified to some $e_{k}$ lifts to a tetrahedron $\widetilde{\Delta}\in\widetilde{\mathbf{\Delta}}^{*}$, then $\{e_{k}\}_{k}$ is in fact a singleton and the sense for every normal quadrilateral in band and branch-point tetrahedra is zero.
There are three edges of a branch-point tetrahedron $\widetilde{p}_{i}$ which are lifts of interior edges of $\mathcal{T}$; if one of these three edges is in $\{e_{k}\}_{k}$, then all three are included say $e_{1},e_{2},e_{3}$. For any normal quadrilateral type in $\widetilde{p}_{i}$ we have the sum of the senses given by $\sum_{k=1}^{3}\epsilon_{k}=0$.
In a band tetrahedron $b_{i}$ there is exactly one edge, say $\langle 01\rangle$, in the material boundary of $\mathcal{T}$. Let $d^{Q}_{1}$ represent the normal quadrilateral type that separates the edges $\langle 01\rangle$ and $\langle 23\rangle$; let $d^{Q}_{2},d^{Q}_{3}$ be the remaining two types. As $S$ is compatible, the number of normal quadrilaterals in $S$ of types $d^{Q}_{2},d^{Q}_{3}$ are $q_{2}=q_{3}=0$. The sign of a normal quadrilateral of type $d^{Q}_{1}$ with respect to the edge $\langle 23\rangle$ is $\delta_{1}=0$. If one of the (unoriented) edges $\langle 02\rangle,\langle 03\rangle,\langle 12\rangle,\langle 13\rangle$ of $\widetilde{b}_{i}$ are in $\{e_{k}\}_{k}$, then all for must be in $\{e_{k}\}_{k}$. Denote these edges by $e_{1},e_{2},e_{3},e_{4}$ respectively (they must all be distinct by the inflation construction). Then for each type $d^{Q}_{i}\in\{d^{Q}_{1},d^{Q}_{2},d^{Q}_{3}\}$ we have $\sum_{k}\epsilon_{k,i}q_{i}=0$. See Figure 8 for an illustration of the signs that contribute to these sense calculations.
@@
Let $B$ be the boundary linking surface in $\mathcal{T}$. For a normal quadrilateral of $B$ in a band tetrahedron $\widetilde{b}_{i}$ there are two opposite edges $\varepsilon_{1}$ and $\varepsilon_{2}$ which crush to an edge in the frame with which $\mathcal{T}$ was inflated (see Figure 3). Through normal isotopies $B$ and $S$ can be positioned such that the intersection $\Gamma$ of $S$ with $B$ has no arcs in the normal quadrilaterals of $B$ that are isotopic to $\varepsilon_{1},\varepsilon_{2}$. Crushing those cells of $\mathcal{C}_{X}$, the cell-decomposition of the complement of $B$ not containing the vertex of $\mathcal{T}$, which are not truncated-tetrahedra yields a cell-decomposition $\bar{\mathcal{T}}$ of $M$ consisting of truncated tetrahedra. By the same arguments of Lemma 3.4 in [6], crushing these cells maps normal discs of $S$ to discs in $\bar{\mathcal{T}}$. As $S$ is compatible, the above Q-matching equations show that these discs glue together to form a surface $\bar{S}$ that is isotopic to $S$ in $M$. Specifically, crushing induces a cell-like map between $\bar{S}$ and $S\cap\mathcal{C}_{X}$. By [3, 38] $\bar{S}$ and $S\cap\mathcal{C}_{X}$ are homeomorphic. By the preceding arguments, $\bar{S}$ is a core of a spun-normal surface, the normal quadrilaterals of which determined by those normal quadrilaterals of $S$ in $\widetilde{\mathbf{\Delta}}^{*}$. Hence $\bar{S}$ is a core of $S^{*}$ completing the proof.
While in general for a band tetrahedron $\widetilde{b}_{i}$ the edge $\langle 23\rangle$ may be identified to the same edge class in $\mathcal{T}$ as one of (unoriented) $\langle 02\rangle,\langle 03\rangle,\langle 12\rangle,\langle 13\rangle$, the ideal triangulations used in Lemma 3.6 are 0-efficient [21] and thus have no faces which are 3-folds or dunce hats [25]. Whence, in this work these three edge classes will be distinct.
As incompatible normal quadrilaterals intersect the material boundary of $\mathcal{T}$, any closed normal surface must be compatible. This gives an alternate proof to Theorem 3.5 of [6] in the case of the inflation triangulation being a short inflation.